01/25/17 – TEI submits SOSS for project. Project pending SOSS approval and PO.
11/17/16 – Hamilton County requested SOSS.
02/16/17 – TEI, Facilities and Building Manager meet on site to kick off project.
02/22/17 – Trane has provided first cut sheet on modular unit for this job, we are trying to fit it into different locations, we are going with a 20 ton unit.
03/01/07 – ThermalTech sent preliminary equipment locations to the County and requested fieldwork to follow up on schematic design. We could potentially have a design meeting at this point and solidify where the unit will be installed.
03/09/17 – ThermalTech updated the floor plans after another field trip down to the site and proposed a final location of one 20-ton unit on the 6th floor with locations for new ductwork and a new wall to enclose the new unit. The county to follow up with Sheriff about using this space for a new unit. TTE moving forward with schedules and Trace runs.
03/15/17 – Revised cost estimates sent to Bert. We are still overbudget but not as badly after electrical modified their cost. We are waiting on revised data from GBBN and THP prior to making additional changes to the estimate. We can make the 6th floor unit work but now its a capital investment issue. Load calcs are complete and we can serve these spaces with a 20 ton, 10,000 cfm unit and we believe we can fit all the adjoining ductwork together to make that happen, the unit would go on the 6th floor.
03/17/17 – TEI contacts GBBN and THP, requesting Architectural and Structural respectively, assistance on this project and included the 5th and 6th floor proposed work drawings for reference.
03/20/17 -50% design dwgs were sent for review to Bert, Tony and Andy. The construction estimate is a little higher than the 192,000 originally budgeted. We are currently trying to finalize the general and structural numbers.
03/20/17 – GBBN sends estimating input to ThermalTech on Architectural implications of the project.
03/21/17 – Spoke to John Milar at THP and we will meet on Thursday to review the AHU locations for the Alternate bids.
03/23/17 – THP – Project explored different approaches not to provide platforms. Meeting now planned to evaluated platforms on 3/23/17.
04/26/16 – Jail Holding unit was added to the design and added as an Alternate to the job. New review drawings were sent to County on 4/25. Additional budgeting was requested for additional design and change orders by owner.
04/27/17 – THP reviewed field conditions and analysis original drawings. Proposed unit and location do not require modification to the structural frame.
05/12/17 – Project is 90% designed but in holding pattern awaiting bid results from the JC Platforms.
06/26/17 – Received revised budgets from Trane on this project and worked on relief air scenario a little bit more. Bid opening could be in July for this project.
06/28/17 – Render the Jury Room for Judge Shannon to see the new ductwork and VAV box.
07/20/17 – GBBN and TEI meet with Facilities for a Field Verification session on site.
07/24/17 – Design team met on site with GBBN to start wall and ceiling work. Some abatement is required at windows we are using for this project. Wall layout confirmed. ThermalTech updated HVAC drawings and they are ready for review. It was decided the Jury room would have an L Soffit instead of a new ceiling, TEI to work with GBBN on this and add access panels for VAV box maintenance.
07/24/17 – GBBN meets with Pinnacle Engineering and Facilities to examine the extent of abatement required around the existing windows and wall peeling paint. Facilities confirms abatement will be included in the Bid Package for this project. Abatement required is the removal of the interior window glazing putty (with window removal), scraping of the peeling and loose lead paint and removal of all pigeon droppings on the sills and wall surfaces. This will be the same for all four (4) windows in this area affected by the scope of work.
07/27/17 – Controls price came in yesterday for $20,000 to upgrade the system for the new layout. This includes removing three units and extending new controls to the new unit named AHU-53N. ThermalTech proposed new smaller soffit and VAV in mezzanine, county okay with this approach.
08/21/17 – Meeting with TEI and GBBN to view area and existing windows to the south of the visitation booths.
08/30/17 – TEI will get ITB and estimate complete this week, project to be Issued for Bid on 9/7/17.
08/31/17 – BW stated that he thinks the drawings are good to go. TTC stated he needed to update the HVAC details for access through the visitation booth add notes for painting the louvers to match the outside of the building color on that row.
09/19/17 – The prebid meeting was conducted today with good turnout by many contractors, some mechanical and some general contractors attended.
09/25/17 – Addendums 1 & 2 have been issued for this project. The bid opening is scheduled for Oct 12 for this project.
10/24/17 – Addendums 3,4,5 have been issued and the Bid Opening extended to Nov 2 for budget issues and equipment clarifications.
11/14/17 – TEI and GBBN attend Post Bid Meeting with Apparent Low Bidder at TEI’s Offices.
11/28/17 – GBBN and ThermalTech have update the drawings to Permit status and removed all the work not being installed and updated the floor plans to make the fifth floor room the new Mechanical Room. Some work remains and some coordination still remains but we believe it will definitely work it is just small spaces and not a lot of free room around the equipment.
02/08/18 – Const Mtg held. Submittals are coming in and being returned to contractor. ThermalTech met Triton on site for fieldwork.
02/19/18 – All HVAC submittals returned. Permit cost worked out with Triton, they can follow up with the City for costs.
02/20/18 – Door, Frame and Door Hardware submittals stamped “Approved as Noted” and returned to ThermalTech.
03/12/18 – Pre-construction meeting held with Triton for this project. Project schedule, SOV, and additional submittals required from Triton. Permit not ready yet, Triton to work out final cost with city, ThermalTech has worked out all the details with the city plans examiner. Triton to distribute meeting notes and revised schedule.
03/21/18 – VAV submittal returned for Tuttle and Bailey with revised btus per TEI markups. Fire Protection & Controls submittals still need review by TEI.
03/24/18 – HVAC Permit is Issued per EZ-trak. Electrical permit is reviewed and on hold for pickup by EC. Next Meeting is Monday. Triton reported today that the room is not clear for construction yet.
04/23/18 – Const Mtg #4 – Triton is on site working on demo and new work. AH-53 has been removed from KP area. New ductwork is being installed. Existing pipe has been turned off and some of the existing piping has been capped off. A leak in Cooper area may be related to our work and shutting down our heating hot water to AH-53. Work is on schedule.
04/09/18 – Const Mtg #3 held and meeting notes distributed. We worked on some of the supply duct routes and return paths.
04/23/18 – Const Mtg #4 held and meeting notes distributed. Reroute return duct KP area, submit CO for sprinkler piping change, finalize HHW and CHW lines to and from units and some of the piping in Shanahan area will be below the duct.
05/07/18 – Const Mtg #5 held and meeting notes distributed. Honeywell on site and starting to work, interior lighting in AHU discussed, const mtgs changed to weekly for final run, Triton to cover all vents in Shanahan’s area and clean debris up daily in work area.
05/14/18 – Const Mtg #6 held and meeting notes distributed. Triton asked to submit change orders for extra work on this project so far. Shanahan soffit will be converted to rigid board ductwork and pants fitting to be fixed before insulation.
05/21/18 – Const MTg #7 held and meeting notes distributed. We are wrapping up Shanahan courtrooms and moving over to Cooper area. Honeywell controls are in place and starting to control the unit. IBI inspection schedule for next week KP ceiling should be going back up, change orders approved, painting has started and we are moving into the final phase of the project.
06/26/18 – Final inspection was approved last Monday. Mechanical and Electrical punchlist have been distributed. Request for Change Order Work #2 has been designed by TEI and sent to Triton for pricing. We have a rough idea of contingency monies left on this project and will try to fashion the changes with that remaining funding.
07/16/18 – Met AAB on site for testing of the airflow through the AHU. We tested the unit at various speeds and took videos on the inner workings of the unit and discovered at 45hz water migrates from the drain pan to the fan section under the drain pan divider. We will contact the manufacturer to rectify this issue.
07/25/18 – Final punchlist was back checked this month. The water leaked returned to which TTE and Triton have been working to resolve with JCI almost daily – there is a lot of behind the scenes work and communications Facilities not included on discussing issues and remedies for this problem. As of yesterday, Triton has increased the depth of the trap, resealed the interior of the unit, and tested interior drainage. JCI tested the levelness of the unit and found it to be installed properly, water was poured into the drain pan and found to drain back and down to the outlet of the unit and at 40hz all seemed to be well.
07/26/18 – The unit was set up to 45Hz and allowed to run overnight at the higher speed. If the unit is still dry we will release it to automatic operation.
08/28/18 – TTC went on site with Triton to work out the reheat coil fitting above the mezzanine in Shanahan’s area. It’s a very tight fit and the two other VAV boxes have to have their supply ducts move so the new zoning works. Other than that one coil, everything else was fitting well. TTE approved controls submittals last week and confirmed with Randy Waddell yesterday for final order.
08/29/18 – Triton reported things are going well but contacted Honeywell to see when they will be on site adding controls and piping to the system.
09/24/18 – ThermalTech met with Triton last week to hash our some piping changes that were not making sense in the field and Triton is making those changes this week, followed by insulator and controls contractor wrapping up. All work to be done by Thursday September 27. Triton still working at night. TEI still looking for hot water reheat lockout temperatures from Honeywell.
10/23/18 – All change order work is complete and the system is programmed into the controls. Space temperatures look good today and the heating appears to be on. All spaces are 72 and above. TEI will start the closeout process on the project.
11/29/18 – Triton dropped off the OM’s yesterday to ThermalTech and TTC hand delivered them to the County including final payapps and closeout paperwork at the Partner Meeting. This project can be archived.
01/22/19 – Project running good. TEI collected all sensors from the space except for the one in the Sheriff surveillance space. Hopefully Barth will recover that for me.
01/23/19 – Project Ready to Archive.
01/25/17 – TEI submits SOSS for project. Project pending SOSS approval and PO.